# Superconducting Cloud Chamber

### v1

@@ Appendix A Appendix A – Detection of the Phase Variation
$\phi=\arctan\left(\frac{-\sum\theta(t)\sin(2\pi f_{\pi}t)}{-\sum\theta(t)\cos(2\pi f_{\pi}t)}\right).$ (7)
Here, $\theta(t)$ represents the discrete sampling of the SQUID response, and $f_{m}=n\times f_{\rm ramp}$ is the modulation frequency. Typically, $f_{m}$ is required to be at least $10$ times the maximum signal frequency.
@@ Appendix B Appendix B – Multiplexing Factor
The number of RF SQUIDs in a readout module, which we define as the multiplexing factor $M$, can be estimated using as follows. Let $S_{f}$ be the frequency spacing between resonator tones, and $f_{\rm ADC}$ be the ADC bandwidth, the multiplexing factor is then defined as $M\equiv f_{\rm ADC}/S_{f}$. Several constraints apply to $S_{f}$ HOLMES:2019ykt : (1) $S_{f}$ must be significantly larger than the resonator bandwidth, $S_{f}>g_{f}\times\Delta f_{\rm BW}$, where $\Delta f_{\rm BW}$ is the resonantor bandwidth and $g_{f}$ is a guard factor; (2) The Nyquist-Sannon sampling theorem requires $\Delta f_{\rm BW}>2f_{m}$; (3) The ramp frequency must exceed the maximum signal frequency, $f_{\rm ramp}>R_{d}/\tau$, where $\tau$ is the rise time of the charged particle (DM) signal and $R_{d}$ is a signal distortion factor. In conclusion, the multiplexing factor per ADC board is,

### v2

@@ Appendix A Appendix A – Detection of the Phase Variation
$\phi=\arctan\left(\frac{-\sum\theta(t)\sin(2\pi f_{\text{mod}}t)}{-\sum\theta(t)\cos(2\pi f_{\text{mod}}t)}\right).$ (8)
Here, $\theta(t)$ represents the discrete sampling of the SQUID response, and $f_{\text{mod}}=n\times f_{\rm ramp}$ is the modulation frequency. Typically, $f_{\text{mod}}$ is required to be at least $10$ times the maximum signal frequency.
@@ Appendix B Appendix B – Multiplexing Factor
The number of RF SQUIDs in a readout module, which we define as the multiplexing factor $M$, can be estimated using as follows. Let $S_{f}$ be the frequency spacing between resonator tones, and $f_{\rm ADC}$ be the ADC bandwidth, the multiplexing factor is then defined as $M\equiv f_{\rm ADC}/S_{f}$. Several constraints apply to $S_{f}$ HOLMES:2019ykt : (1) $S_{f}$ must be significantly larger than the resonator bandwidth, $S_{f}>g_{f}\times\Delta f_{\rm BW}$, where $\Delta f_{\rm BW}$ is the resonantor bandwidth and $g_{f}$ is a guard factor; (2) The Nyquist-Sannon sampling theorem requires $\Delta f_{\rm BW}>2f_{\text{mod}}$; (3) The ramp frequency must exceed the maximum signal frequency, $f_{\rm ramp}>R_{d}/\tau$, where $\tau$ is the rise time of the charged particle (DM) signal and $R_{d}$ is a signal distortion factor. In conclusion, the multiplexing factor per ADC board is,
@@
Appendix C Appendix C – Determination of Particle Nature
In the main text, we mentioned that by analyzing the signal amplitudes and timing sequences from multiple RF-SQUIDs in the superconducting cloud chamber, we can reconstruct a particle’s trajectory, charge, and velocity. In this appendix, we demonstrate how to extract this information by analyzing signals from any four arbitrarily selected JJs within a localized region of the chamber.
Figure 5: A schematic illustration of a particle passing through the superconducting cloud chamber (local view).
As shown in Fig. 5, the charged particle flies by the detector $1\sim 4$ with distance $d_{i}(i=1\sim 4)$. Correspondingly, the detectors see the phase difference $\phi_{i}(i=1\sim 4)$ and the signal peak arises at the moment $t_{i}(i=1\sim 4)$. According to Eq. (2), we can list the following four equations,
$\phi_{i}=\frac{\epsilon e^{2}}{\pi}\frac{\delta}{v_{\chi}d_{i}}.$ (10)
However, the particle charge $\epsilon$ and velocity $v_{\chi}$ are degenerate information that cannot be directly determined from the equations above. At this stage, by introducing the assumption that the particle’s trajectory forms an angle $\theta$ with the detector array, we can establish the relationship between the measured distances,
$d_{3}=d_{1}+\frac{D}{\cos\theta}\,,\quad d_{4}=d_{2}+\frac{D}{\cos\theta}.$ (11)
We can also establish the relationship between the signal timing and the particle’s velocity,
$t_{2}-t_{1}=\frac{D}{\cos\theta v_{\chi}}.$ (12)
We have now formulated a system of seven independent equations, enabling the determination of all seven unknowns, $\epsilon$, $v_{\chi}$, $\theta$, and $d_{i}(i=1\sim 4)$. From them, we can figure out the charge and velocity of the flying particle.
To see the time difference, the timing resolution between different detectors must be better than, $\delta t<t_{2}-t_{1}\simeq D/v_{\chi}$. Since $D\sim 1\,$mm and $v_{\chi}\subset(0.01,1000)\,$m/s, the time resolution should be better than $\delta t<10^{-5}\,$s. This condition is easily achieved by the current timing technology.